A four-year project requires an initial investment of $208,000 for fixed assets and $19,500 for net working capital. All of the net working capital will be recouped at the end of the project. The annual operating cash flow is $76,200 and the discount rate is 6.00 percent.

I. Create a timeline/chart to identify the relevant Cash Flows and their timing

II. Use the Cash Flows for the following analysis:


a) What is the Payback Period? b) What is the NPV? c) What is the IRR? d) What is the Profitability Index? e) Should the company do this project? Why or Why not?
Answer #1
Year 0 1 2 3 4
Cash flow -2,27,500 76,200 76,200 76,200 76,200
Cumulative cash flow formula =76200 =76200+76200 =76200+76200+76200 =76200+76200+76200+76200
Cumulative cash flow 76,200 1,52,400 2,28,600 3,04,800
As cumulative cash flow exceeds the initial investment and working capital in year 3
Payback period = 3 years
NPV $                            36,541.05
IRR 12.82%
Profitability Index 1.16
Company should do this project as it has positive NPV and IRR higher than the discount rate